OpenStudy (anonymous):

Anyone want a Medal? Which of the following is equal to the square root of the cube root of 6? 6 to the power of 1 over 6 6 to the power of 1 over 3 6 to the power of 2 over 3 6 to the power of 3 over 2

OpenStudy (jango_in_dtown):

cube root of 6 =\[6^{1/3}\]

OpenStudy (jango_in_dtown):

and its square root =\[(6^{1/3})^{1/2}\]

OpenStudy (jango_in_dtown):

now (a^m)^n=a^mn

OpenStudy (jango_in_dtown):

so the solution is 6^(1/3*1/2)=6^(1/6)

OpenStudy (anonymous):

Thanks!